OpenStudy (anonymous):

The net force acting on a car rolling down a ramp is the addition of three forces . One of the forces is the ramp pushing up to support the car . Name the other two forces acting on the car . And two opposing forces on the car . Can someone please tell me what they are and why ?

OpenStudy (anonymous):

If it is a real car (not a car toy) moving downward, I can think in theses forces: One of the forces is the ramp pushing up to support the car: This I think is the Normal Force \(N\), but since \(Ncos(\theta) \neq w\), the car slides down with a force equal to the component of the weight parallel to the slope: \(wsin(\theta)\). If there is friction \(f_s\), it will oppose the downward motion. Anpther force can be the force exerted by the engine. So you have the following (assuming it is referring to the net force making the car move downward with the presence of friction): \(wsin(\theta) + F_{engine} = f_s\)

OpenStudy (anonymous):

If the forces in the left-hand side are greater than the force of friction, the car's wheels will start to slip.

OpenStudy (surry99):

friction and gravity